# Semiparametric Skew-Elliptical Distributions For High-Dimensional Graphical Models

### v1

@@ 2.2 The Nonparanormal SKEPTIC
Both $\hat{\rho}_{ij}$ and $\hat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$. Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, respectively. The cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$ are denoted by $F_{i}$ and $F_{j}$. The population versions of Spearman’s rho and Kendall’s tau are defined as $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively. For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$. We recall the SKEPTIC estimator.
@@
(Kruskal, (1958), Liu et al., 2012a ): Assuming $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$. From these results, we define two estimators for the unknown correlation matrix $(\Sigma^{0}_{ij})$ that are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$,
Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable. This concept will be fundamental in the definition of the elliptical skew-(S)KEPTIC estimator.
2.3 The meta-elliptical distribution
The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in the same way as the nonparanormal of Liu et al., (2009) do for Gaussian distributions. The meta-elliptical family is a strict extension of the nonparanormal distribution.
@@ 2.3 The meta-elliptical distribution
[…]
@@
(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is defined as the latent generalized correlation matrix.
For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix. Nevertheless, the meaning of a missing edge is different with respect to the nonparanormal case. Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence. On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means conditional linear independence.
The meta-elliptical family contains many useful distributions among which asymmetric distributions. However, the estimate of the correlation matrix does not directly take into account the skewness parameters. We extend the meta-elliptical family giving a direct contribution of the skewness parameters for the conditional independence relations, and thus considering the $\mathcal{E}$ family of distributions.
2.4 $\mathcal{E}$ distributions
We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in Arellano-Valle and Genton, 2010b . If a random vector has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, where $h^{(p+q)}$ is the density generator. We focus on two specifications of the family which are: the unified skew-$t$ ($\operatorname{SUT}$) distribution (Arellano-Valle and Genton, 2010a ), and the unified skew-Normal ($\operatorname{SUN}$) distribution (Arellano-Valle and Azzalini, (2006)). Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into $\operatorname{SUN}$ and $\operatorname{SUT}$ distributions. Formally, we define the $\operatorname{SU\mathcal{E}}$ distribtuion.
Definition 2.4.
(Arellano-Valle and Genton, 2010b ) A continuous p-dimensional random vector $X$ has a multivariate unified skew-elliptical ($\mathcal{E}$) distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
$f_{X}(x)=\frac{1}{F_{q}\left(\tau;\Gamma+\Lambda\bar{\Sigma}\Lambda^{T},h^{(q)}\right)}f_{p}\left(y;\mu,\Sigma,h^{(p)}\right)F_{q}\left(\Lambda z+\tau;\Gamma,h_{Q(z)}^{(q)}\right),$ (4)
where $w=\sigma^{-1}(y-\mu),Q(w)=w^{\top}(\Sigma^{0})^{-1}w,\Lambda$ is a $q\times p$ real matrix controlling shape, $\tau\in\mathbb{R}^{q}$ is the extension parameter, $\Gamma$ is a $q\times q$ positive definite correlation matrix, $f_{p}\left(y;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(w))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$, positive definite $p\times p$ dispersion matrix $\Sigma$, with $p\times p$ scale and correlation matrices $\sigma=\operatorname{diag}(\Sigma)^{1/2}$ and $\Sigma^{0}=\sigma^{-1}\Sigma\sigma^{-1}$, respectively, and density generator $h^{(p)},F_{r}\left(y;\Sigma,h^{(r)}\right)$ denotes the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(q)}$, and $h_{Q(w)}^{(q)}(u)=h^{(p+q)}(u+Q(w))/h^{(p)}(Q(w))$.
$\operatorname{SU\mathcal{E}}$ $h^{(p)}$
$\operatorname{SUN}$ $h^{(p)}=\phi^{(p)}$$(u)$where: $\phi^{(p)}(u)=\frac{1}{(2\pi)^{p/2}}\exp\left(-\frac{u^{2}}{2}\right)$
$\operatorname{SUT}$ $h^{(p)}=t^{(p)}(u,\nu)$ where: $t^{(p)}(u,\nu)=\frac{\Gamma\left(\frac{\nu+p}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\left(\pi a\right)^{p/2}}\left(1+\frac{u}{\nu}\right)^{-\frac{\nu+p}{2}}$
The relation between the $\operatorname{SUN}$ and the $\operatorname{SU\mathcal{E}}$ is defined through the following stochastic representation.
Proposition 2.5.
(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, when the extension parameter $\tau=0$, we can write $\boldsymbol{X}$ as a mixture of a $\operatorname{SUN}$ distribution: $\boldsymbol{X}=\mu+\sigma V^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim SUN_{p,q}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
The $\operatorname{SUN}$ distribution results in a different parameterization of the closed skew-Normal (CSN) distribution in Domìnguez-Molina et al., (2003). The CSN distribution corresponds to the case in which the cumulative distribution functions in Definition 2.4 are the Gaussian cumulative distribution functions. Therefore, the $\operatorname{SU\mathcal{E}}$ distribution can be expressed in terms of the $\operatorname{CSN}$ distribution. This is useful for deriving the moments and for sampling random numbers from the $\operatorname{SU\mathcal{E}}$ distribution. The $\operatorname{CSN}$ distribution has the following stochastic representation as a combination of a Normal distribution and a half-Normal distribution.
Definition 2.6.
(Domìnguez-Molina et al., (2003)) Let $\boldsymbol{Y}\sim\operatorname{CSN}_{p,q}(\mu,\Sigma,\Gamma,0,\Delta)$, the random vector $\boldsymbol{Y}$ admits the following stochastic representation:
[…]
where: $\boldsymbol{U}\sim\operatorname{N}_{p}(0,I_{p})$ and $\boldsymbol{V}\sim\operatorname{HN}_{q}(0,\Delta+\Gamma\Sigma\Gamma)$, the Normal and the half-Normal distribution, respectively.
The stochastic representation in (5) allows us to estimate an undirected decomposable graph.
Definition 2.7.
(Zareifard et al., (2016)) Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(\mu,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
$f_{\boldsymbol{X}}(x)=2^{p}\phi_{p}\left(x;\mu,\Omega^{-1}\right)\Phi_{p}\left(D_{\alpha}L(x-\mu);0,D^{-1}_{\kappa}\right),$ (6)
where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(\mu,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(\mu,\Omega^{-1})$.
We introduce an estimator for skew-Gaussian graphical models defined in Zareifard et al., (2016) that extends the SKEPTIC including a skewness component to accommodate non-symmetric data.
3 The elliptical skew-(S)KEPTIC estimator
Definition 3.1.
@@ 3 The elliptical skew-(S)KEPTIC estimator
(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ and a nonnegative random variable $\xi$, such that
$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU}\mathcal{E}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma),$ (7)
@@
We exploit the stochastic representation of $\operatorname{SU\mathcal{E}}$ distribution through a CSN distribution (Proposition 2.5) to define the skew-elliptical graphical model. That is, we propose a nonparametric rank-based estimator, which we call elliptical skew-(S)KEPTIC estimator, based on the stochastic representation (5), according to the parametrization of $\operatorname{CSN}$ distribution in (6):
$\boldsymbol{Z}\stackrel{{\scriptstyle p}}{{=}}\boldsymbol{\mu}+L^{-1}D_{\kappa}^{-\frac{1}{2}}D_{\alpha}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{U}+L^{-1}D_{\kappa}^{-\frac{1}{2}}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{V},$ (8)
where: $\boldsymbol{U}\sim N_{p}(0,I_{p})$ and $\boldsymbol{V}\sim HN_{p}(0,I_{p})$, where $N_{p}(0,I_{p})$ is a $p$-dimensional Normal distribution, and $HN_{p}(0,I_{p})$ is a $p$-dimensional half-Normal distribution. The distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent. The first two moments of $\boldsymbol{Z}$ are:
$\mathbb{E}(\boldsymbol{Z})$ $=\mu+L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$

### v2

@@ 2.3 The meta-elliptical distribution
We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in Arellano-Valle and Genton, 2010b . If a random vector $\boldsymbol{X}$ has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, where $h^{(2p)}$ is the density generator. We focus on two specifications of the family, which are: the closed skew-normal (CSN) distribution of Domìnguez-Molina et al., (2003) [^1] We build on the closed skew-normal (CSN) distribution, which represents a reparametrization of the unified skew-normal (SUN) distribution employed by Arellano-Valle and Genton, 2010b . and the unified skew-$t$ (SUT) distribution (Arellano-Valle and Genton, 2010a ). Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into CSN and SUT distributions. Formally, we define the $\operatorname{SU\mathcal{E}}$ distribtuion.
Definition 2.4.
(Arellano-Valle and Genton, 2010b, ) A continuous p-dimensional random vector $\boldsymbol{X}$ has a multivariate unified skew-elliptical $\operatorname{(SU\mathcal{E})}$ distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
$f_{X}(x)=2^{p}f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)F_{p}\left(\Lambda z;\Gamma,h_{Q(z)}^{(p)}\right),$ (4)
where $\mu$ is the location parameter, $\Sigma$ is a $p\times p$ dispersion matrix, $h^{(p)}$ is the density generator, $F_{r}\left(x;\Sigma,h^{(r)}\right)$ the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(r)}$. Additionally, $z=\sigma^{-1}(x-\mu)$, $\Lambda$ is a $p\times p$ real matrix controlling shape, $Q(z)=z^{\top}(\Sigma^{0})^{-1}z$, $\Gamma$ is a $p\times p$ positive definite symmetric matrix, $f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(z))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$. In conclusion, $h_{Q(z)}^{(p)}(u)=h^{(2p)}(u+Q(z))/h^{(p)}(Q(z))$.
$\operatorname{SU\mathcal{E}}$ $h^{(p)}$
CSN $h^{(p)}=\phi^{(p)}$$(u)$  where: $\phi^{(p)}(u)=\frac{1}{(2\pi)^{p/2}}\exp\left(-\frac{u^{2}}{2}\right)$
SUT $h^{(p)}=t^{(p)}(u,\nu)$ where: $t^{(p)}(u,\nu)=\frac{\Gamma\left(\frac{\nu+p}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\left(\pi a\right)^{p/2}}\left(1+\frac{u}{\nu}\right)^{-\frac{\nu+p}{2}}$
The relation between the CSN and the $\operatorname{SU\mathcal{E}}$ is defined through the following stochastic representation.
Proposition 2.5.
(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, we can write $\boldsymbol{X}$ as a mixture of a CSN distribution: $\boldsymbol{X}=\mu+\sigma V_{0}^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
@@
(Arellano-Valle and Genton, 2010b, ) If  $\text{ $V_{0}$}\sim Gamma(\nu/2,\nu/2)$, then
$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\mu,\Sigma,\Lambda,\nu,0,\Gamma)$.
We construct an undirected graph on the closed skew-normal distribution, following Zareifard et al., (2016).
@@
Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
$f_{\boldsymbol{X}}(x)=2^{p}\phi_{p}\left(x-\mu;0,\Omega^{-1}\right)\Phi_{p}\left(D_{\alpha}L(x-\mu);0,D^{-1}_{\kappa}\right),$ (5)
where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, in other words: $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$. Further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(0,\Omega^{-1})$.
The graphical model for the unified skew-elliptical distribution can be constructed by utilizing the sparsity structure of the closed skew-normal distribution. We rely on a semiparametric approach that considers the class of unified skew-elliptical distributions to extend the SKEPTIC estimator of Liu et al., 2012a proposing a robust estimator of graphical models that accounts for skewness and heavy tails.
3 The elliptical skew-(S)KEPTIC estimators
@@
(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ such that
$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma),$ (6)
@@
We employ the stochastic representation of the unified skew-elliptical distribution, characterized by a closed skew-normal distribution $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ (Proposition 2.5), to introduce the elliptical skew-(S)KEPTIC estimators for constructing skew-elliptical graphical models. The closed skew-normal distribution in (5) can be expressed in the following stochastic form:
$\boldsymbol{X}\stackrel{{\scriptstyle d}}{{=}}L^{-1}D_{\kappa}^{-\frac{1}{2}}D_{\alpha}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{U}+L^{-1}D_{\kappa}^{-\frac{1}{2}}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{V},$ (7)
where: $\boldsymbol{U}\sim HN_{p}(0,I_{p})$ and $\boldsymbol{V}\sim N_{p}(0,I_{p})$, where $HN_{p}(0,I_{p})$ is a $p$-dimensional half-normal distribution, and $N_{p}(0,I_{p})$ is a $p$-dimensional normal distribution. The distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent. The first two moments of $\boldsymbol{X}$ are:
${E}(\boldsymbol{X})$ $=L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$
$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}D_{\kappa}^{\frac{1}{2}}(I-D^{2})D_{\kappa}^{\frac{1}{2}}L\right)^{-1},$ (8)
where $D=\sqrt{\frac{2}{\pi}}D_{\alpha}(I_{p}+D^{2}_{\alpha})^{-\frac{1}{2}}$.
Theorem 3.4 (elliptical skew-KEPTIC estimator).
Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$. The elliptical skew-KEPTIC estimator is:
As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with the Pearson correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanormal. The proof of the Theorem is in Appendix A.
Lemma 3.5.
(elliptical skew-SKEPTIC estimator) Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution. When $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$ reduces to the CSN distribution in (5), the elliptical skew-SKEPTIC estimator is:
We generally refer to the estimators in Theorem 3.4 and Lemma 3.5 as elliptical skew-(S)KEPTIC estimators. To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimators, we need to estimate the parameters of the CSN distribution. The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^2] The estimated parameters come from the skew-normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..
4 Methods for the Estimation of the Precision Matrix and the Undirected Graph
4.1 The Dantzig selector
The Dantzig selector estimator (Yuan,, 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix. A sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$. The Dantzig procedure involves two steps:
  • Estimation: for $j=1,\ldots,p$, calculate